# Neural Combinatorial Optimization Algorithms for Solving Vehicle Routing Problems: A Comprehensive Survey with Perspectives

### v2

@@ V Learning to Predict (L2P) Solvers
Figure 6: Illustration of the information prediction processes of L2P-O and L2P-M solvers. L2P-O solvers aim to enhance OR algorithms by providing valuable information solely before the searches begin. Subsequently, OR algorithms address VRP instances by leveraging the predicted information along with problem definitions. Conversely, L2P-M solvers collaborate with OR algorithms continuously during the search processes, offering prediction of key information at each decision step. These L2P-M solvers take states of the OR algorithms, encompassing problem definitions, as their inputs [47].
@@ V-A Learning to Predict-Once (L2P-O) Solvers
A group of L2P-O solvers predict information only once before commencing searches with OR algorithms, aiming to expedite the search processes. For example, to confine the search space and thus expedite the search processes of DP, Kool et al. [44] adopted a Graph Convolution Networks (GCN) to predict a sparse heatmap for each TSP instance, which highlights promising edges. Subsequently, DP excludes solutions with fewer “heated” edges, effectively eliminating a multitude of inappropriate solutions and confining the search space, thereby reducing runtime. Similar to [44], other studies [151, 152, 153] also employ L2P-O solvers to predict the scores of edges, effectively confining the search space and accelerating the search processes. In addition, it is worth mentioning that certain studies [37, 154, 155, 68] adopted search algorithms, e.g., Monte Carlo Tree Search (MCTS) and Beam Search, rather than OR algorithms, to construct solutions based on the predicted heatmaps. Nevertheless, because the solution construction process of these methods employing search algorithms (e.g., [37]) is closely similar to OR algorithms (e.g., [44]), we categorize these methods as L2P-O solvers as well. For example, Joshi et al. [37] leveraged GCN to predict heatmaps and employed Beam Search on the resulting heatmaps to construct solutions. Specifically, they formulated TSP instances as 2D Euclidean graphs and utilized the optimal solutions as labels for training GCN to predict heatmaps. While constructing solutions using search algorithms is more efficient compared to the OR algorithms, it comes with the inadequacy of being unable to solve multi-constrained VRPs.
@@
In summary, certain L2P-O solvers exploit the predicted information to expedite the search processes of OR algorithm, while others leverage such information to enhance the quality of solutions generated by OR algorithms. Meanwhile, in addition to the advantages aforementioned, it should not be overlooked that L2P-O solvers also face the generalization capability challenge, similar to L2C and L2I solvers.
@@ V-B Learning to Predict-Multiplicity (L2P-M) Solvers
Concorde (2007) [144$\dagger$ OR 5.696 - 9m 7.765 - 43m - - - - - -
HGS (2022) [145$\dagger$ OR - - - - - - 10.366 - 1.2d 15.563 - 1.5d
GCN (2019) [37$\dagger$ L2P-O/BS 5.698 0.04% 23m 7.869 1.35% 46m - - - - - -
Att-GCN (2021) [163$\dagger$ L2P-O/MCTS 5.691 0.01% 8m 7.764 0.04% 15m - - - - - -
CAVE-Opt-DE (2021) [46$\dagger$ L2P-M/DE - 0.02% 2.5d - 0.34% 1.8d 10.400 - 4.7d 15.750 - 11d
@@
To strive for high-quality solutions, Hottung et al. [46] exploited an NN to predict the latent search space with a smooth landscape, then employed OR algorithms (e.g., Differential Evolution (DE) [169]) to search in the predicted space. Recently, Liu et al. [170] proposed a framework based on Large Language Models (LLMs), to automatically generate multiple heuristic algorithms for producing high-quality solutions for VRPs. Similar to Evolutionary Computation (EC), the proposed LLM framework encompasses initialization, selection, crossover, and mutation steps. However, unlike EC, where each individual represents a solution of a VRP instance, in the LLM framework, each individual embodies a heuristic algorithm [170]. In addition, during each iteration, crossover and mutation operators involve employing prompts (comprising task descriptions, parent individuals, and generation instructions) to guide LLMs in generating multiple heuristic algorithms. Following [170], Ye et al. [166] further proposed two innovative prompt methods for crossover and mutation operators, respectively. The short-term prompt is employed for the crossover operator, compelling LLMs to analyze the differences in performance between the two parents. Conversely, the long-term prompt is utilized for the mutation operator, prompting LLMs to assimilate insights from historical individuals. While these LLM-generated heuristic algorithms exhibit strong generalization capabilities, their effectiveness in addressing large-scale VRPs remains a concern. Encouraging LLMs to generate more complex heuristic algorithms could potentially address this challenge.
@@ VI Inadequacies, Efforts, and Perspectives of Existing NCO Solvers
Despite the high-level performance achieved by existing NCO solvers in terms of the quality of solutions and time complexity, they still face certain inadequacies when compared to the conventional OR algorithms [18, 171, 172, 173, 174]. Sections VI-A to VI-D introduce four specific inadequacies of existing NCO solvers and propose corresponding directions aimed at addressing these inadequacies, respectively.
@@ VI-A Inadequacies of Generalization Capability
The prevailing inadequacy among the majority of existing NCO solvers is their limited capacity for generalization across diverse data distribution patterns [175, 176, 177, 178]. For example, solvers trained with instances following the uniform distribution (i.e., $\forall v_{i}\in\bm{V},v_{i}\sim\mathcal{U}([0,1]^{2})$) are less likely capable of handling instances conforming to the normal distribution (i.e., $\forall v_{i}\in\bm{V},v_{i}\sim\mathcal{N}([0,1]^{2})$) [58]. This inadequacy impedes the application of NCO solvers in real-world problems [23].
@@
In addition, NCO solvers exhibit limited generalization capability across different problem scales simultaneously [57]. For example, those trained on trivially small scales struggle to generalize to larger instances. While the approaches introduced in the previous paragraph solely focus on distribution, recent studies [182] and [183] simultaneously address generalization concerning both the problem scale and distribution. Specifically, Manchanda et al. [182] formalized solving VRP instances over a given instance distribution and scale as a separate learning task. Subsequently, they utilized Meta Learning to learn a model capable of adapting to various tasks. Following [182], Zhou et al. [183] replaced the random task sampling method used in [182] with a hierarchical scheduler to adaptively select different types of distribution. In addition, to mitigate the computing overhead, they introduced a second-order derivative approximation method within the Meta Learning framework. This approach entails utilizing second-order derivatives primarily in the early stages of training, and subsequently relying exclusively on their first-order counterparts.
@@
While considerable advancement has been made in enhancing the generalization of NCO solvers, it is crucial to acknowledge that these generalization methods require more hyperparameters and efforts on data synthesis, resulting in a significant overhead in computing resources. We propose employing a lightweight NN to adjust the attention scores of Transformer-based solvers for different problem scales. This approach can enhance generalization capability while requiring fewer computing resources. In addition, these solvers fall under the purview of domain adaptation for addressing the out-of-distribution problem [184]. Specifically, these solvers focus on adapting to the instances from known target distributions. However, they exhibit inadequacies in effectively handling instances from unknown distribution patterns in real-world problems (e.g., TSPLib [185] and CVRPLib [186]). We suggest incorporating Invariant Learning into NCO solvers, directing their attention to data features unaffected by environmental variations, thus enhancing their generalization capability. Finally, drawing from the emphasis on enhancing the generalization capacity of L2C solvers in prior studies, attention to other solver categories, such as L2I, L2P-O, and L2P-M solvers, has been relatively scarce. We posit that with the help of OR algorithms, L2P-O and L2P-M solvers could potentially demonstrate superior generalization capabilities.
@@ VI-B Inadequacies of Solving Large-scale VRPs
where $\exp(\cdot)$ and $\top$ denote exponential and transpose functions, respectively. Vectors $q_{i}$, $k_{i}$, and $v_{i}$ denote the query, key, and value of the $i$th node, respectively. According to (15), it is evident that each instance involves a dot product computation with a complexity of $O(N^{2}d)$ and a memory usage complexity of $O(N^{2})$ [131]. When the problem scale $N$ increases to $\lambda N$, where $\lambda$ denotes an integer greater than one, the complexity of dot product computation and memory usage grow to $O(\lambda^{2}N^{2}d)$ and $O(\lambda^{2}N^{2})$, respectively. Consequently, as the problem scale expands, solvers employing the self-attention mechanism are prone to memory depletion during training, hampering convergence [36]. Similarly, various other NNs experience problem scale-related effects to varying degrees.
To address this issue, certain methods have embraced the D&C manner to decompose large-scale problems into multiple small-scale sub-problems [62, 163, 189, 148, 80]. Subsequently, these methods employ diverse solvers to construct or improve solutions for all decomposed sub-problems, followed by the integration of all sub-problems. For example, Pan et al. [62] proposed H-TSP, which utilizes an upper-level partition model and a lower-level L2C solver for partitioning and solving sub-problems, respectively. Similarly, Hou et al. [80] exploited a Transformer to divide a problem into multiple sub-problems and incorporated LKH3 to construct solutions for each sub-problem. For these D&C solvers, identifying the optimal decomposition strategy for large-scale VRPs stands as a crucial factor [148].
@@
While pursuing the same objective, other NCO solvers [68, 128, 135, 75] diverge from the D&C manner in addressing the time and memory expenses associated with solving large-scale VRPs. Specifically, they adopt lighter NNs or innovative training paradigms. For example, Yang et al. [135] replaced scaled dot-product attention with sampled scaled dot-production attention (see Section III-A for more details). In addition, to optimize memory usage, Luo et al. [36] adopted SL as the training paradigm to train an NN rather than RL. This choice stems from the fact that RL training necessitates generating the complete solution before computing the reward, implying a substantial demand for memory and computing resources. Furthermore, in contrast to the original study [37], which involves training a GCN using SL, Min et al. [68] introduced an innovative UL loss function to train GCN. By adopting this design, GCN utilizes approximately 10% of the number of parameters compared to the SL training paradigm.
@@
Table VI presents the performance of NCO solvers designed for solving large-scale TSP and CVRP instances. For TSP, comparing L2C solvers that abstain from employing the D&C approach (TSPformer [135] and Pointerformer [128]) with their counterparts utilizing this strategy (H-TSP [62]) exposes a consistent performance gap ranging from 1% to 10%. However, even though H-TSP adopts the D&C manner, its capability is limited to handling TSP instances with a maximum of 1,000 nodes. In contrast, L2P-O solvers outperform L2C solvers, especially the recently proposed heatmap-based solvers (e.g., UTSP [68], DIFUSCO [45], and T2TCO [165]) that can achieve high-level performance in real-time, with gaps falling within the 1% range, when dealing with 500-node and 1,000-node TSP instances. Notably, on the challenging 10,000-node TSP, DIFUSCO [45] achieved an impressive gap value of 2.58%, credited to the training method exploiting the diffusion model. Specifically, DIFUSCO constantly adds Bernoulli noises [190] to optimal solutions of instances and compels a GCN to denoise the added noises. Through this training method, GCN has the capability to predict high-quality heatmaps. Subsequently, MCTS is employed to construct (sub-)optimal solutions based on these high-quality heatmaps. For CVRP, current NCO solvers (e.g., TAM-LKH3 [80]) adopt the D&C manner and utilize LKH3 to solve decomposed sub-problems, resulting in enhanced performance that surpasses that of LKH3 alone.
@@
However, all these heatmap-based solvers lack the generality to efficiently handle VRP variants beyond TSP [42], e.g., CVRP. Because the intricate constraints of VRP variants cannot be adequately modeled solely through the predicted heatmaps, as mentioned in Section V-C. We propose to employ more sophisticated OR algorithms (e.g., DP) for heatmap-based solvers, rather than relying solely on basic search algorithms. The incorporation of OR algorithms presents a viable direction to this problem. In addition, we suggest utilizing more lightweight NNs (e.g., Mamba [191]) as replacements for the widely adopted Transformer in L2C solvers. This design adjustment enables NCO solvers to solve VRPs from a global perspective, thereby mitigating the distance cost associated with inadequate decomposition strategies in the D&C manner.

### v3

@@ V Learning to Predict (L2P) Solvers
Figure 7: Illustration of the information prediction processes of L2P-O and L2P-M solvers. L2P-O solvers aim to enhance OR algorithms by providing valuable information solely before the searches begin. Subsequently, OR algorithms solve VRP instances by leveraging the predicted information along with problem definitions. Conversely, L2P-M solvers collaborate with OR algorithms continuously during the search processes, offering prediction of key information at each decision step. These L2P-M solvers take states of the OR algorithms, encompassing problem definitions, as their inputs [50].
@@ V-A Learning to Predict-Once (L2P-O) Solvers
A group of L2P-O solvers predict information only once before commencing searches with OR algorithms, aiming to expedite the search processes. For example, to confine the search space and thus expedite the search processes of DP, Kool et al. [47] adopted a Graph Convolution Networks (GCN) to predict a sparse heatmap for each TSP instance, which highlights promising edges. Subsequently, DP excludes solutions with fewer “heated” edges, effectively eliminating a multitude of inappropriate solutions and confining the search space, thereby reducing runtime. Similar to [47], other studies [161, 162, 163] also employed L2P-O solvers to predict the scores of edges, effectively confining the search space and accelerating the search processes. In addition, it is worth mentioning that certain studies [40, 164, 165, 73] adopted search algorithms, e.g., Monte Carlo Tree Search (MCTS) and Beam Search, rather than OR algorithms, to construct solutions based on the predicted heatmaps. Nevertheless, because the solution construction process of these methods employing search algorithms (e.g., [40]) is closely similar to OR algorithms (e.g., [47]), we categorize these methods as L2P-O solvers as well. For example, Joshi et al. [40] leveraged GCN to predict heatmaps and employed Beam Search on the resulting heatmaps to construct solutions. Specifically, they formulated TSP instances as 2D Euclidean graphs and utilized the optimal solutions as labels for training GCN to predict heatmaps. While constructing solutions using search algorithms is more efficient compared to the OR algorithms, it comes with the inadequacy of being unable to solve multi-constrained VRPs.
@@
In summary, certain L2P-O solvers exploit the predicted information to expedite the search processes of OR algorithms, while others leverage such information to enhance the quality of solutions generated by OR algorithms. Meanwhile, in addition to the advantages aforementioned, it should not be overlooked that L2P-O solvers also face the generalization capability inadequacy, similar to L2C and L2I solvers.
@@ V-B Learning to Predict-Multiplicity (L2P-M) Solvers
Concorde (2007) [151$\dagger$ OR 5.696 - 9m 7.765 - 43m - - - - - -
HGS (2022) [152$\dagger$ OR - - - - - - 10.366 - 1.2d 15.563 - 1.5d
GCN (2019) [40$\dagger$ L2P-O/BS 5.698 0.04% 23m 7.869 1.35% 46m - - - - - -
Att-GCN (2021) [173$\dagger$ L2P-O/MCTS 5.691 0.01% 8m 7.764 0.04% 15m - - - - - -
CAVE-Opt-DE (2021) [49$\dagger$ L2P-M/DE - 0.02% 2.5d - 0.34% 1.8d 10.400 - 4.7d 15.750 - 11d
@@
To strive for high-quality solutions, Hottung et al. [49] exploited an NN to predict the latent search space with a smooth landscape, then employed OR algorithms (e.g., Differential Evolution (DE) [180]) to search in the predicted space. Recently, Liu et al. [181] proposed a framework based on Large Language Models (LLMs), to automatically generate multiple heuristic algorithms for producing high-quality solutions for VRPs. Similar to Evolutionary Computation (EC), the proposed LLM framework encompasses initialization, selection, crossover, and mutation steps. However, unlike EC, where each individual represents a solution of a VRP instance, in the LLM framework, each individual embodies a heuristic algorithm [181]. In addition, during each iteration, crossover and mutation operators involve employing prompts (comprising task descriptions, parent individuals, and generation instructions) to guide LLMs in generating multiple heuristic algorithms. Following [181], Ye et al. [176] further proposed two innovative prompt methods for crossover and mutation operators, respectively. The short-term prompt is employed for the crossover operator, compelling LLMs to analyze the differences in performance between the two parents as search directions. Conversely, the long-term prompt is utilized for the mutation operator, prompting LLMs to assimilate insights from historical individuals. To enable LLMs to provide specific search directions, Wu et al. [177] proposed the Core Abstraction Prompting method, which abstracts the core components from elite heuristics and incorporates them as prior knowledge in short- and long-term prompts. Future research could further encourage LLMs to develop more sophisticated heuristic algorithms that may yield superior performance. Moreover, a prior study [182] suggested that heuristics can enhance the performance of NCO solvers. This also represents a promising direction for improving the generalization ability of NCO solvers across different distributions and scales by deriving tailored heuristics through LLMs.
@@ VI Inadequacies, Efforts, and Perspectives of Existing NCO Solvers
Despite the high-level performance achieved by existing NCO solvers in terms of the quality of solutions and time complexity, they still face certain inadequacies when compared to the conventional OR algorithms [21, 183, 184, 185, 186]. Sections VI-A to VI-D introduce four specific inadequacies of existing NCO solvers, discuss existing efforts to solve these inadequacies, identify open inadequacies, and propose additional promising directions to address them.
@@ VI-A Inadequacies of Generalization Capability
The prevailing inadequacy among the majority of existing NCO solvers is their limited capacity for generalization across diverse data distribution patterns [187, 188, 189, 190, 191, 192]. For example, solvers trained with instances following the uniform distribution (i.e., $\forall v_{i}\in\bm{V},v_{i}\sim\mathcal{U}([0,1]^{2})$) are less likely capable of handling instances conforming to the normal distribution (i.e., $\forall v_{i}\in\bm{V},v_{i}\sim\mathcal{N}([0,1]^{2})$) [61]. This inadequacy impedes the application of NCO solvers in real-world problems [27].
@@
In addition, NCO solvers exhibit limited generalization capability across different problem scales simultaneously [60]. For example, those trained on trivially small scales struggle to generalize to larger instances. While the approaches introduced in the previous paragraph solely focus on distribution, certain studies [196, 197] simultaneously addressed generalization concerning both the problem scale and distribution. Specifically, Manchanda et al. [196] formalized solving VRP instances over a given instance distribution and scale as a separate learning task. Subsequently, they utilized Meta Learning to learn a model capable of adapting to various tasks. Following [196], Zhou et al. [197] replaced the random task sampling method used in [196] with a hierarchical scheduler to adaptively select different types of task. In addition, to mitigate the computing overhead, they introduced a second-order derivative approximation method within the Meta Learning framework. This approach entails utilizing second-order derivatives primarily in the early stages of training, and subsequently relying exclusively on their first-order counterparts.
@@
While considerable advancement has been made in enhancing the generalization of NCO solvers, it is crucial to acknowledge that these generalization methods require more hyperparameters and efforts on data synthesis, resulting in a significant overhead in computing resources. We propose employing more lightweight NNs to improve generalization capability. For example, Xiao et al. [198] proposed a plug-and-play entropy-based scaling factor to adjust attention scores, thereby enhancing the generalization capability of solver while requiring fewer computing resources. In addition, these solvers fall under the purview of domain adaptation for addressing the out-of-distribution problem [199]. Specifically, these solvers focus on adapting to the instances from known target distributions. However, they exhibit inadequacies in effectively handling instances from unknown distribution patterns in real-world problems (e.g., TSPLib [200] and CVRPLib [201]). We suggest incorporating Invariant Learning into NCO solvers, directing their attention to data features unaffected by environmental variations, thus enhancing their generalization capability. Recent studies [63, 59, 186] emphasized the role of nearest-neighbor relationships among nodes in capturing local invariance across problems of varying scales and distributions.
Currently, certain solvers achieve high-level performance on synthetic TSP datasets across different distributions and scales. For instance, the SOTA solver GELD [59] attains gap values of 2.38%, 3.44%, 4.7%, and 2.81% on 10,000-node TSP instances with uniform, clustered, explosion, and implosion distributions, respectively. However, on real-world instances of the same scale, such as those in TSPLib [^2] URL: http://comopt.ifi.uni-heidelberg.de/software/TSPLIB95/ and National TSPs [^3] URL: https://www.math.uwaterloo.ca/tsp/world/countries.html, GELD’s gap value remains high at 7% due to differences between the real-world distributions and the training distributions. This highlights the need for further improvement of NCO solvers on real-world instances. In addition, drawing from the emphasis on enhancing the generalization capacity of L2C solvers in prior studies, attention to other solver categories, such as L2I, L2P-O, and L2P-M solvers, has been relatively scarce. We posit that with the help of OR algorithms, L2P-O and L2P-M solvers could potentially demonstrate superior generalization capabilities.
@@ VI-B Inadequacies of Solving Large-scale VRPs
where $\exp(\cdot)$ and $\top$ denote exponential and transpose functions, respectively. Vectors $q_{i}$, $k_{i}$, and $v_{i}$ denote the query, key, and value of the $i$th node, respectively. According to (15), it is evident that each instance involves a dot product computation with a complexity of $\mathcal{O}(N^{2}d)$ and a memory usage complexity of $\mathcal{O}(N^{2})$ [141]. When the problem scale $N$ increases to $\lambda N$, where $\lambda$ denotes an integer greater than one, the complexity of dot product computation and memory usage grow to $\mathcal{O}(\lambda^{2}N^{2}d)$ and $\mathcal{O}(\lambda^{2}N^{2})$, respectively. Consequently, as the problem scale expands, solvers employing the self-attention mechanism are prone to memory depletion during training, hampering convergence [39]. Similarly, various other NNs experience problem scale-related effects to varying degrees.
@@
To address this issue, certain methods have embraced the D&C manner to decompose large-scale problems into multiple small-scale sub-problems [65, 173, 204, 157, 85]. Subsequently, these methods employ diverse solvers to construct or improve solutions for all decomposed sub-problems, followed by the integration of all sub-problems. For example, Pan et al. [65] proposed H-TSP, which utilizes an upper-level partition model and a lower-level L2C solver for partitioning and solving sub-problems, respectively. Similarly, Hou et al. [85] exploited a Transformer to divide a problem into multiple sub-problems and incorporated LKH3 to construct solutions for each sub-problem. For these D&C solvers, identifying the optimal decomposition strategy for large-scale VRPs stands as a crucial factor [71, 157].
@@
While pursuing the same objective, other NCO solvers [73, 136, 144, 206] diverge from the D&C manner in addressing the time and memory expenses associated with solving large-scale VRPs. Specifically, they adopt lighter NNs or innovative training paradigms. For example, Yang et al. [144] replaced scaled dot-product attention with sampled scaled dot-production attention (see Section III-A for more details). Xiao et al. [59] proposed the Region Average Linear Attention mechanism, which partitions nodes into regions to facilitate efficient global information exchange, achieving a computation complexity $\mathcal{O}(Nd)$. In addition, to optimize memory usage, Luo et al. [39] adopted SL as the training paradigm to train an NN rather than RL. This choice stems from the fact that RL training necessitates generating the complete solution before computing the reward, implying a substantial demand for memory and computing resources. More recently, studies [207] and [208] exploited the high-quality solutions produced by the solver as labels, enabling self-supervised learning. Furthermore, in contrast to the original study [40], which involves training a GCN using SL, Min et al. [73] introduced an innovative UL loss function to train GCN. By adopting this design, GCN utilizes approximately 10% of the number of parameters compared to the SL training paradigm.
@@
Table VI presents the performance of NCO solvers designed for solving large-scale TSP and CVRP instances. For TSP, comparing L2C solvers that abstain from employing the D&C approach (TSPformer [144] and Pointerformer [136]) with their counterparts utilizing this strategy (H-TSP [65]) exposes a consistent performance gap ranging from 1% to 10%. However, even though H-TSP adopts the D&C manner, its capability is limited to handling TSP instances with a maximum of 1,000 nodes. In contrast, L2P-O solvers outperform L2C solvers, especially the recently proposed heatmap-based solvers (e.g., UTSP [73], DIFUSCO [48], T2TCO [175], and Fast T2T [209]) that can achieve high-level performance in real-time, with gaps falling within the 1% range, when dealing with 500-node and 1,000-node TSP instances. Notably, on the challenging 10,000-node TSP, DIFUSCO [48] achieved an impressive gap value of 2.58%, credited to the training method exploiting the diffusion model. Specifically, DIFUSCO constantly adds Bernoulli noises to optimal solutions of instances and compels a GCN to denoise the added noises. Through this training method, GCN has the capability to predict high-quality heatmaps. Subsequently, MCTS is employed to construct (sub-)optimal solutions based on these high-quality heatmaps. To accelerate inference, Wang et al. [205] proposed DEITSP, which optimizes multiple generation steps of DIFUSCO into a single step, transforming Bernoulli noise into high-quality heatmaps through a specially designed bimodal graph Transformer for one-step prediction.
@@
For CVRP, a VRP variant beyond TSP, heatmap-based solvers lack the generality to handle it efficiently [45, 205]. Because the intricate capacity constraint of CVRP cannot be adequately modeled solely through the predicted heatmaps, as mentioned in Section V-C. We propose to employ more sophisticated OR algorithms (e.g., DP) for heatmap-based solvers, rather than relying solely on basic search algorithms. The incorporation of OR algorithms presents a viable direction to this problem. To solve large-scale CVRP instances, current NCO solvers generally adopt the D&C manner. For example, Zheng et al. [71] proposed a unified neural D&C framework (UDC), which employs an Anisotropic Graph Neural Network for global node partitioning and subsequently adopts an L2C solver to solve all CVRP sub-problems, achieving performance that surpasses LKH3.
Although certain solvers have attained strong performance on synthetic large-scale TSP and CVRP instances, their effectiveness on other VRP variants still requires improvement. For instance, although UDC [71] employs the same L2C solver (i.e., ICAM [206]) for both CVRP and Open VRP (OVRP), its performance on the 2,000-node OVRP instances lags behind LKH3 by 8.71%. For future work in this regard, we suggest utilizing more lightweight NNs (e.g., Mamba [210]) as replacements for the widely adopted Transformer in L2C solvers. This design adjustment enables L2C solvers to solve VRP with constraints from a global perspective, thereby mitigating the distance cost associated with inadequate decomposition strategies in the D&C manner. In addition, as mentioned in Section VI-A, there is a performance gap between existing solvers on synthetic datasets and real-world ones, which is an inadequacy that warrants further investigation in the future.